2 oranges and 5 apples cost $2.50 3 oranges and 4 apples cost $2.70 find the cost of an orange and an apple
Answer by Paul(988)

LEt the cost of the organges be x
Let the cost of apples be y
Now:
2x+5y=2.50 --> SOlve for x
2x=2.50-5y
x=1.25-2.5y (subsitution)

Now:
3x+4y=2.70
Subsitute for x
3(1.25-2.5y)+4y=2.70
3.75-7.5y+4y=2.70
-3.5y=-1.05
y=0.30
x=1.25-2.5(0.30)
x=0.50

Hence a cost of an organge is $0.50 and a cost of an apple is $0.30.
Paul.